OpenStudy (anonymous):

College Algebra: reviewing for Final Exam! If the number 3 is a zero of a polynomial function f(x), then which of the following is true? a.) x-3 is a factor of f(x) b.) x+3 is a factor of f(x) c.) -3 is a zero of f(x) d.) f(-3)=0 Can someone help me answer this question and understand polynomials?

OpenStudy (bahrom7893):

I'm not going to go into too much details regarding polynomial functions, basically they're some functions the equation of which looks like this:\[f(x)=a*x^n+b*x^{n-1}+c*x^{n-2}....+some \space constant*x+some \space constant\]

OpenStudy (p0sitr0n):

x-3 is a factor

OpenStudy (bahrom7893):

where all the powers are positive integers

OpenStudy (p0sitr0n):

By the fundamental theorem of algebra i believe, you can factor (x-r) out of any polynomial that has root r

OpenStudy (fibonaccichick666):

You would have to know there is a root of 3 though

OpenStudy (fibonaccichick666):

oh, and we do haha

OpenStudy (p0sitr0n):

it is stated

OpenStudy (bahrom7893):

Now, that may look a bit overwhelming, but here are a few simple examples:\[f(x) = x+1\]\[f(x)=x \space (which \space is \space the \space same \space as \space x+0)\]

OpenStudy (fibonaccichick666):

^Note, that has a zero at x=-1

OpenStudy (bahrom7893):

But stop giving the answers, I'm trying to explain here

OpenStudy (anonymous):

I'm reading @bahrom7893 don't worry. I'm reading each one of your comments. But I'm learning. Trying to comprehend what you're saying.

OpenStudy (bahrom7893):

Another thing special about polynomials are their graphs.. you're probably very familiar with this one:\[y=x^2\]